how many calories are required to convert 100 grams of ice to steam?
1
Expert's answer
2020-02-05T12:13:56-0500
"Q_{ice}=100(80)=8000\ cal"

"Q_{water}=100(1)(100)=10000\ cal"

"Q_{steam}=100(540)=54000\ cal"

"Q=8000+10000+54000=72000\ cal"